OpenStudy (anonymous):

(1+x)(1-x)(1-x+x^2)(1+x+x^2) = ? (1+x)(1-x)(1-x+x^2)(1+x+x^2) = ? @Mathematics

OpenStudy (anonymous):

\[(a) 1+x^6\] \[(b) 1+x^3\] \[(c) 1-x^6\] \[(d) 1+x^4\]

OpenStudy (anonymous):

1-x^6

OpenStudy (anonymous):

How?

OpenStudy (ash2326):

(1+x)(1-x+x^2)(1-x)(1-x+x^2) (1+x^3)(1-x^3) 1-x^6

OpenStudy (anonymous):

(1+x)(1-x+x^2)=1+x^3 (1-x)(1+x+x^2)=1-x^3 (1+x^3)(1-x^3)=1-x^6

OpenStudy (anonymous):

Thanks to both of you.. : )